OpenStudy (anonymous):

Differentiate the function K(x)=2^tanx

OpenStudy (anonymous):

2 to the power of tanx? This seems like a weird function but I suggest you use the chain rule:\[\frac{ da }{ db }=\frac{ da }{ dc } \times \frac{ dc }{ db }\]

OpenStudy (anonymous):

Let m(x) = 2^x and n(x) = tan(x) Thus find the derivative of m(n(x)), w/c is m'(n(x)) * n"(x) Therefore the final answer will be\[\ln(2){\times}2^{\tan(x)}\sec^{2}(x)\]

OpenStudy (anonymous):

Yup, that seems right. Incidentally, you can check your answer by typing the function into: http://www.wolframalpha.com/ It can differentiate things for you.